OpenStudy (anonymous):

A 60 foot light pole casts a shadow that is 40 feet long. At the same time, a nearby light post casts a shadow that is 15 feet long. How tall is the light post?

OpenStudy (anonymous):

tan(2/3)*15

OpenStudy (anonymous):

Yay, proportions! 60/40=x/15 x=900/40 The light post is 22.5 feet tall.